### v1

# A General Latent Embedding Approach for Modeling Non-uniform High-dimensional Sparse Hypergraphs with Multiplicity

Abstract
Recent research has shown growing interest in modeling hypergraphs, which capture polyadic interactions among entities beyond traditional dyadic relations. However, most existing methodologies for hypergraphs face significant limitations, including their heavy reliance on uniformity restrictions for hyperlink orders and their inability to account for repeated observations of identical hyperlinks. In this work, we introduce a novel and general latent embedding approach that addresses these challenges through the integration of latent embeddings, vertex degree heterogeneity parameters, and an order-adjusting parameter. Theoretically, we investigate the identifiability conditions for the latent embeddings and associated parameters, and we establish the convergence rates of their estimators along with asymptotic distributions. Computationally, we employ a projected gradient ascent algorithm for parameter estimation. Comprehensive simulation studies demonstrate the effectiveness of the algorithm and validate the theoretical findings. Moreover, an application to a co-citation hypergraph illustrates the advantages of the proposed method.
Keywords: hypergraphs, identifiability conditions, maximum likelihood estimation, entry-wise consistency, asymptotic distribution.
1 Introduction
Network analysis explores relations among entities. Most existing network analysis tools focus on analyzing dyadic relations, i.e., those between two entities, through the examination of graphs (Newman, 2018). However, real-life relations often go beyond pairwise interactions. For instance, in co-authorship data (Ji & Jin, 2016; Ji et al., 2022), researchers typically form groups larger than pairs when co-authoring papers. In co-sponsorship data (Fowler, 2006), legislators endorsing legislation as co-sponsors form clusters that usually involve more than two people. In gene interaction data (Razick et al., 2008), subsets of genes, rather than just pairs, associate to potentially form protein complexes. Modeling and analyzing these higher-order relations, as highlighted by Ke et al. (2019), yield deeper insights that are often overlooked by traditional graph analysis. These higher-order interactions can be naturally captured and aggregated by hypergraphs, which allows for a more comprehensive analysis.
We denote a hypergraph by ${\cal H}({\cal V}_{n},{\cal E}_{m})$, where ${\cal V}_{n}=\{v_{1},v_{2},\cdots,v_{n}\}$ represents the collection of $n$ vertices (individuals) and ${\cal E}_{m}=\{e_{1},e_{2},\cdots,e_{m}\}$ represents the collection of $m$ hyperlinks/hyper-edges (interactions). For notation simplicity, we denote $[n]=\{1,\cdots,n\}$ and ${\cal V}_{n}=[n]$. Each hyperlink $e_{j}$, for $j\in[m]$, consists of a subset of vertices that form the $j$-th hyperlink; that is, $e_{j}\subseteq[n]$ with $i\in e_{j}$ if and only if vertex $i$ is part of the $j$-th hyperlink. For example, if vertices $1,2$ and $3$ form the first hyperlink, we have $e_{1}=\{1,2,3\}$. A hyperlink consisting of $d$ vertices is termed an order-$d$ hyperlink. If all hyperlinks $e_{j}\in{\cal E}_{m}$ have the same order $d$, the hypergraph ${\cal H}({\cal V}_{n},{\cal E}_{m})$ is termed a $d$-uniform hypergraph. For instance, a typical graph network is essentially a 2-uniform hypergraph without repetitive hyperlinks. Conversely, if the hyperlinks in ${\cal E}_{m}$ vary in their orders, the hypergraph is described as non-uniform. Non-uniform hypergraphs are ubiquitous in the real world. For example, consider co-authorship relationships, where authors are vertices and each paper is treated as a hyperlink that includes its co-authors. The number of co-authors for each paper, i.e., the order of the hyperlink, naturally varies across different papers. Furthermore, it is common for the same set of co-authors to collaborate on multiple papers, leading to identical hyperlinks appearing multiple times within a hypergraph. This phenomenon is referred to as the multiplicity of hyperlinks. The non-uniformity and multiplicity of hyperlinks pose unique challenges in modeling hypergraphs.
In this work, we introduce a latent embedding framework to model general hypergraphs that accommodates the non-uniformity and multiplicity of hyperlinks under a high-dimensional setup where both $n$ and $m$ go to infinity. Assume that each vertex is associated with a latent embedding vector (or vertex embedding), denoted by $\{\mathbf{z}_{i}\}_{i\in[n]}\subset\mathbb{R}^{K}$, where $K$ represents the dimension of the latent embedding space (vertex space). Additionally, each hyperlink also carries a corresponding latent embedding (or hyperlink embedding), denoted by $\{\mathbf{f}_{j}\}_{j\in[m]}\subset\mathbb{R}^{K}$, potentially residing in another latent space (the hyperlink space). Conditioned on $\{\mathbf{z}_{i}\}_{i\in[n]}$ and $\{\mathbf{f}_{j}\}_{j\in[m]}$, our model assumes that the appearances of each node on the hyperlinks are independent. In other words, interactions among vertices are captured by these latent embeddings. Such conditional independence, given (hyper)graph embeddings, has been considered in many (hyper)graph models, including the stochastic block model (Holland et al., 1983), the latent space network model (Hoff et al., 2002), and the hypergraph degree-corrected block model (Ke et al., 2019). Specifically, consider a hypothetical hypergraph ${\cal H}([n],\{E_{1},\ldots,E_{m}\})$, where each $E_{j}$ is a random subset of $[n]$. The observed hypergraph ${\cal H}([n],\{e_{1},\ldots,e_{m}\})$ is a realization of ${\cal H}([n],\{E_{1},\ldots,E_{m}\})$. Given the latent embeddings, we assume that $\{E_{1},\ldots,E_{m}\}$ are independent with the following distribution:
$\mathbb{P}(E_{j}={e}|\boldsymbol{\Xi}_{m,n})=\prod_{i\in{e}}p_{ji}\prod_{i\in[n]\setminus{e}}(1-p_{ji}),~{}~{}~{}\text{for any }{e}\subseteq[n],$ (1.1)
where $p_{ji}=g\big{(}\beta_{m,n}+\alpha_{i}+{\cal K}(\mathbf{f}_{j},\mathbf{z}_{i})\big{)}\in$ $(0,1)$ and $\boldsymbol{\Xi}_{m,n}=\big{\{}\beta_{m,n},\{\alpha_{i}\}_{i=1}^{n},\{\mathbf{z}_{i}\}_{i=1}^{n},\{\mathbf{f}_{j}\}_{j=1}^{m}\big{\}}$. Here, ${\cal K}(\cdot,\cdot):\mathbb{R}^{K}\times\mathbb{R}^{K}\to\mathbb{R}$ measures the “similarity” (to be defined) between the two latent embeddings, and $g(\cdot):\mathbb{R}\to[0,1]$, a monotonically increasing function, maps a real value to a probability. The $\{\alpha_{i}\}_{i\in[n]}$ parameters (with $\alpha_{i}\in\mathbb{R}$) account for heterogeneity in vertex degrees, which can be interpreted as a popularity measure of the vertices, and $\beta_{m,n}\in\mathbb{R}$, the order-adjusting parameter, controls the overall order of hyperlinks. Unlike most existing hypergraph modeling approaches where the order of hyperlinks is uniform, our model accommodates varying hyperlink orders. However, the orders of hyperlinks typically remain much smaller than both the number of vertices within the hypergraph and the number of observed hyperlinks, which poses an additional modeling challenge. $\beta_{m,n}$ serves to address this challenge by monitoring the overall order of hyperlinks compared to $n$ and $m$. In model (1.1), $\{\mathbf{z}_{i}\}_{i\in[n]}$ and $\{\mathbf{f}_{j}\}_{j\in[m]}$ can be treated either as realizations of some embedding distribution or as parameters to estimate. We treat them as fixed parameters in this work. In situations where these embeddings are treated as random vectors, our analysis can be regarded as making statistical inferences conditional on them. Moreover, it is worth noting that different $E_{j}$’s have different distributions. This means that rather than assuming an identical generating mechanism for all $m$ hyperlinks, our framework permits individualized probability assignments for each observed hyperlink, which is more flexible in modeling real-world hypergraphs. More discussion and interpretation of (1.1) are in Section 2.
1.1 Related work
Modeling general hypergraphs has been recognized as both an important task and a significant challenge. A commonly employed strategy is to project the hypergraph onto a weighted graph, enabling analysis using methods such as modularity maximization (Kumar et al., 2018), graph-likelihood-based approaches (Lee et al., 2020), and spectrum-based techniques (Ghoshdastidar & Dukkipati, 2015) on the converted graph. These approaches have been applied in the studies of academic collaborations (Newman, 2001; Ji & Jin, 2016) and congress voting relations (Fowler, 2006; Lee et al., 2017). For instance, in a co-authorship hypergraph, a hyperlink consists of all co-authors of a paper, usually more than two. Transforming the hypergraph into a weighted graph involves considering each pair of authors and constructing pairwise links with weights proportional to the number of papers they co-authored. However, as Ke et al. (2019) pointed out, such projection process results in information loss and may yield sub-optimal performance when applied to community detection tasks.
Recent research has seen a surge in the direct modeling and analysis of hypergraphs. For instance, Ke et al. (2019) proposed a regularized tensor power iteration algorithm called Tensor-SCORE for community detection on uniform hypergraphs. Extensions to non-uniform hypergraphs were suggested without further investigation. The hypergraph degree-corrected block model introduced by Ke et al. (2019) incorporates vertex degree heterogeneity into the framework of the hypergraph planted partition model (Ghoshdastidar & Dukkipati, 2017). Yuan & Qu (2021) developed a tensor-based joint network embedding method that encodes both pairwise links and a subgroup of uniform hyperlinks into a latent space, capturing the dependency between pairwise and multi-way links in predicting potential unobserved hyperlinks. Yuan et al. (2022) conducted an extensive study on hypergraph testing for uniform hypergraphs under a spectrum of hyperlink probability scenarios, providing comprehensive theoretical analysis and significant insights into the subject. Zhen & Wang (2022) introduced a null vertex to augment a non-uniform hypergraph into a uniform hypergraph for community detection purposes. However, most existing hypergraph modeling approaches have significant limitations, including a heavy reliance on uniform hypergraph structures and the inability to account for repeated observations of identical hyperlinks due to the use of adjacency tensor representation. In the co-authorship data example, this means that only combinations of authors of a fixed size can be effectively modeled, and information about the number of papers a group of authors has co-authored is lost. In reality, the number of co-authors can vary for different papers, and the same set of co-authors can write more than one paper. Furthermore, the handling of non-uniform hypergraphs remains ad hoc. There is a pressing demand for a general modeling framework.
1.2 Contributions
Our contribution in this article can be summarized in two main aspects. First, we introduce a novel and general approach that addresses the limitations of most existing methods in hypergraph modeling. Our approach does not require the stringent uniformity restriction on hyperlinks that is often presumed in existing methods, and it accounts for repeated observations of identical hyperlinks, which are overlooked in many current approaches. These advantages distinctly set our methodology apart from existing hypergraph modeling approaches.
Second, we investigate the identifiability conditions for the embeddings and associated parameters, and establish convergence rates for their estimators along with their asymptotic distributions. In Section 3.1, we introduce estimators that achieve $F$-consistency for the “${\boldsymbol{\Theta}}$” parameter (to be defined), as well as for the vertex popularity parameters and the order-adjusting parameter. In Section 3.2, we address the challenge of estimating individual components in $\boldsymbol{\Xi}_{m,n}$. Additional identifiability conditions are necessary to disentangle individual parameters, and we design a penalty term based on these conditions. Notably, this penalty term is not for shrinkage or regularization purposes, but for the implementation of a Lagrangian method to address identifiability issues. Furthermore, we allow the order-adjusting parameter $\beta_{m,n}$ to diverge to $\infty$- ∞ to accommodate the sparsity of the order of hyperlinks. This divergence results in irregular behavior of the likelihood function. We dedicate substantial effort to understanding this effect. Given that $\beta_{m,n}$ is designed to control the rate of hyperlink orders in the model, which will be further illustrated in Section 2, this exploration offers valuable insights into the types of datasets that are well-suited to our methodology. As a result, we establish entry-wise consistent estimators for individual parameters in $\boldsymbol{\Xi}_{m,n}$ and characterize their asymptotic distributions. These theoretical results subsequently permit estimation and inference on the probability terms $p_{ji}$’s.
1.3 Organization of the paper and notations
The remainder of this article is organized as follows. Section 2 provides a detailed introduction and interpretation of the order-adjusted latent embedding approach. Section 3 investigates the identifiability conditions and estimation methods for the parameters and presents the corresponding statistical theory. Section 4 presents numerical results on synthetic hypergraphs. Section 5 applies our proposed methodology to a co-citation hypergraph, which leads to interesting discoveries. Section 6 concludes the paper with a discussion. The supplementary material includes proofs of the main results, technical lemmas and their proofs, and additional numerical results.
Throughout this work, we use regular letters, bold regular letters, and bold capital letters to denote scalars, vectors, and matrices, respectively. For any $a,b\in\mathbb{R}$, we denote their maximum and minimum by $a\vee b=\max(a,b)$ and $a\wedge b=\min(a,b)$, respectively. Given any two sequences $(a_{m,n}),(b_{m,n})$ valued in $\mathbb{R}$, we say $a_{m,n}\lesssim b_{m,n}$ (or $a_{m,n}=O(b_{m,n})$, $b_{m,n}\gtrsim a_{m,n}$) if there exists a universal constant $C>0$ such that $|a_{m,n}|\leq C|b_{m,n}|$. We say $a_{m,n}\asymp b_{m,n}$ if $a_{m,n}\lesssim b_{m,n}$ and $a_{m,n}\gtrsim b_{m,n}$. We write $a_{m,n}\ll b_{m,n}$ (or $a_{m,n}=o(b_{m,n})$) if $\lim_{m,n\to\infty}a_{m,n}/b_{m,n}=0$ and $a_{m,n}\gg b_{m,n}$ (or $a_{m,n}=\Omega(b_{m,n})$) if $\lim_{m,n\to\infty}a_{m,n}/b_{m,n}=\infty$. We use $O_{p}(\cdot)$, $\Omega_{p}(\cdot)$, and $o_{p}(\cdot)$ as the stochastic versions of $O(\cdot)$, $\Omega(\cdot)$, and $o(\cdot)$ for convergence in probability. For any positive integer $a$, we use $[a]$ to denote the set $\{1,2,\ldots,a\}$. For any column vector $\mathbf{a}$ and matrix $\mathbf{A}$, we use $\mathbf{a}^{\top}$ and $\mathbf{A}^{\top}$ to denote their transposes, respectively. For an event $\mathfrak{E}$, the indicator function $\mathbf{1}_{\mathfrak{E}}=1$ if $\mathfrak{E}$ occurs and 00 otherwise. For a vector $\mathbf{a}=(a_{1},\cdots,a_{n})^{\top}$, we use $\|\mathbf{a}\|_{2}$ to denote the standard Euclidean norm, $\|\mathbf{a}\|_{0}=\sum_{i\in[n]}\mathbf{1}_{\{a_{i}\neq 0\}}$ to denote the $L_{0}$ norm (the number of non-zeros), and $\|\mathbf{a}\|_{\infty}=\max_{i}|a_{i}|$ to denote the infinity/maximum/supreme norm. For a matrix $\mathbf{A}=(a_{ji})_{m\times n}$, we use $\|\mathbf{A}\|_{F}$ and $\|\mathbf{A}\|_{2}$ to denote its Frobenius and spectral norms, respectively. We use $\|\mathbf{A}\|_{\max}=\max_{j\in[m],i\in[n]}|a_{ji}|$ to denote the matrix maximum/supreme norm. For a square matrix $\mathbf{A}$, we use $\rho_{\max}(\mathbf{A})$ and $\rho_{\min}(\mathbf{A})$ to denote its maximal and minimal eigenvalues, respectively.
2 The order adjusted latent embedding approach
In this paper, we consider ${\cal K}(\cdot,\cdot)$ in (1.1) as the Euclidean inner product in a $K$-dimensional space and employ the logistic function as $g$. Specifically, for $\mathbf{a},\mathbf{b}\in\mathbb{R}^{K}$, ${\cal K}(\mathbf{a},\mathbf{b})=\mathbf{a}^{\top}\mathbf{b}=\sum_{k=1}^{K}a_{k}b_{k}$, and for $a\in\mathbb{R}$, $g(a)=\sigma(a):=1/\big{(}1+\exp(-a)\big{)}$. Then we have
$p_{ji}=\sigma(\beta_{m,n}+\alpha_{i}+\mathbf{f}_{j}^{\top}\mathbf{z}_{i}).$ (2.1)
We show in this section that using the Euclidean inner product in our model offers a clear and coherent interpretation. Although we focus on the logistic link function in this article, it is worth noting that our theoretical analysis is adaptable and can be effectively applied to a variety of other link functions that exhibit similar characteristics. In contrast to conventional approaches that often focus on modeling the latent embeddings of vertices, with hyperlinks being generated from these vertex embeddings, our method jointly models latent embeddings of vertices and hyperlinks. This joint modeling framework offers more flexibility and a more comprehensive understanding of the hyperlink embeddings.
One concern with the model (2.1) is the identifiability of the parameters, primarily because the probability value in (1.1) is determined by $p_{ji}$, which remains invariant under certain transformations involving $\beta_{m,n},\alpha_{i},\mathbf{f}_{j}$ and $\mathbf{z}_{i}$. To identify $\beta_{m,n}$ and $\alpha_{i}$’s, we impose the constraints $\sum_{i=1}^{n}\alpha_{i}=0$ and $\sum_{j=1}^{m}\mathbf{f}_{j}=\mathbf{0}$. The rationale for these restrictions is derived from the right-hand side of (2.1). If we add a constant to all $\alpha_{i}$’s and simultaneously subtract it from $\beta_{m,n}$, the resulting probability distribution remains unchanged. Similarly, adding an arbitrary vector $\boldsymbol{\mu}_{\mathbf{f}}\in\mathbb{R}^{K}$ to each $\mathbf{f}_{j}$ while adjusting each $\alpha_{i}$ by subtracting $\boldsymbol{\mu}_{\mathbf{f}}^{\top}\mathbf{z}_{i}$ also preserves the original probability distribution. We defer the discussion on the identification of $\mathbf{f}_{j}$’s and $\mathbf{z}_{i}$’s to Section 3.2, where we aim to estimate individual parameters in $\boldsymbol{\Xi}_{m,n}$. Throughout this article, we assume that
$\max\big{\{}\max_{i\in[n]}|\alpha_{i}|,\max_{j\in[m]}\|\mathbf{f}_{j}\|_{2},\max_{i\in[n]}\|\mathbf{z}_{i}\|_{2}\big{\}}\leq C$ (2.2)
for some constant $C>0$. We allow $\beta_{m,n}$ to diverge to $\infty$- ∞ as $(m,n)\to\infty$, thus $p_{ji}$’s can be arbitrarily close to 00 in the asymptotic regime. Consequently, the expected order of hyperlinks (i.e. the expected number of vertices in a hyperlink) can be sublinear with respect to the size of the hypergraph. This is consistent with the observation in practice that the order of hyperlinks often increases at a much slower rate compared to the growth of the number of vertices.
To further understand (2.1), consider a specific hyperlink indexed by $j\in[m]$. For a subset of vertices $e\subset[n]$ with $|{e}|=d$, given $\boldsymbol{\Xi}_{m,n}$, model (1.1) along with (2.1) specifies the conditional probability that these vertices constitute hyperlink $j$ as
$\mathbb{P}\big{(}E_{j}=e|\boldsymbol{\Xi}_{m,n}\big{)}$ (2.3)
$=$ $\prod_{i\in e}\sigma(\beta_{m,n}+\alpha_{i}+\mathbf{f}_{j}^{\top}\mathbf{z}_{i})\prod_{i\in[n]\setminus e}\big{\{}1-\sigma(\beta_{m,n}+\alpha_{i}+\mathbf{f}_{j}^{\top}\mathbf{z}_{i})\big{\}}$
$=$ $\prod_{i\in{e}}\frac{\exp(\beta_{m,n}+\alpha_{i}+\mathbf{f}_{j}^{\top}\mathbf{z}_{i})}{1+\exp(\beta_{m,n}+\alpha_{i}+\mathbf{f}_{j}^{\top}\mathbf{z}_{i})}\cdot\prod_{i\in[n]\setminus{e}}\frac{1}{1+\exp(\beta_{m,n}+\alpha_{i}+\mathbf{f}_{j}^{\top}\mathbf{z}_{i})}$
$=$ $\frac{\prod_{i\in{e}}\exp(\beta_{m,n}+\alpha_{i}+\mathbf{f}_{j}^{\top}\mathbf{z}_{i})}{\prod_{i\in[n]}\big{\{}1+\exp(\beta_{m,n}+\alpha_{i}+\mathbf{f}_{j}^{\top}\mathbf{z}_{i})\big{\}}}$
$=$ $\frac{\big{[}\exp\big{\{}\beta_{m,n}+\sum_{i\in{e}}\alpha_{i}/d+\mathbf{f}_{j}^{\top}(\sum_{i\in{e}}\mathbf{z}_{i}/d)\big{\}}\big{]}^{d}}{\prod_{i\in[n]}\big{\{}1+\exp(\beta_{m,n}+\alpha_{i}+\mathbf{f}_{j}^{\top}\mathbf{z}_{i})\big{\}}}.$
The denominator of this probability remains constant regardless of the specific subset of vertices ${e}$. The numerator depends on a combination of factors: the order-adjusting parameter $\beta_{m,n}$, the vertices’ average popularity $\sum_{i\in{e}}\alpha_{i}/d$, the inner product between the hyperlink’s embedding $\mathbf{f}_{j}$ and the centroid of the vertices’ embeddings $\sum_{i\in{e}}\mathbf{z}_{i}/d$, and the order of the hyperlink $d$. The multiplicative operations in the second line of (2.3) suggest a form of conditional independence among the vertices given the embeddings. Such conditional independence has been widely considered in (hyper)graph network models, for example, by Holland et al. (1983), Hoff et al. (2002), Ke et al. (2019), and Ma et al. (2020). In this work, we do not model the distributions of embeddings as they can be general. Below,
Refer to caption Refer to caption
(a) (b)
Figure 1: An example of the latent embedding vectors when $K=2$. Figures (a) and (b) display the same latent embedding vectors but consider different combinations of vertices on hyperlinks. The sizes of the dots correspond to different vertex degree parameters: the larger the dot, the higher its corresponding $\alpha$ value.
with the hyperlink embedding vector $\mathbf{f}_{j}$ being fixed, we use Figure 1 to illustrate how the vertex embeddings and associated parameters affect the probability term (2.3):
  • Positioning of vertices’ centroid: Figure 1 (a) highlights two groups of vertices, each with their respective centers denoted as $\boldsymbol{\mu}_{\mathbf{z},1}$ and $\boldsymbol{\mu}_{\mathbf{z},2}$. The Euclidean inner product between $\boldsymbol{\mu}_{\mathbf{z},1}$ and $\mathbf{f}_{j}$ is larger than that between $\boldsymbol{\mu}_{\mathbf{z},2}$ and $\mathbf{f}_{j}$. Focusing only on the term $\mathbf{f}^{\top}_{j}(\sum_{i\in{e}}\mathbf{z}_{i}/d)$, the group of vertices centered around $\boldsymbol{\mu}_{\mathbf{z},1}$ tends to have a higher probability of forming hyperlink $j$ than the other group.
  • Average popularity of vertices: In Figure 1 (b), $\mathbf{f}_{j}^{\top}\boldsymbol{\mu}_{\mathbf{z},1}$ and $\mathbf{f}_{j}^{\top}\boldsymbol{\mu}_{\mathbf{z},3}$ are close, and the two groups have the same number of vertices. We consider the term $\sum_{i\in{e}}\alpha_{i}/d$, which we refer to as the average popularity of vertices in ${e}$. Note that larger dot sizes correspond to higher $\alpha$ values in Figure 1. Due to their higher average popularity, the group of vertices centered around $\boldsymbol{\mu}_{\mathbf{z},1}$ has a higher probability of forming hyperlink $j$ than the group centered around $\boldsymbol{\mu}_{\mathbf{z},3}$.
The order-adjusting parameter $\beta_{m,n}$ is not visualized in Figure 1, but it is particularly crucial for determining the orders of hyperlinks. In the considered sparse scenario, where $\beta_{m,n}\to-\infty$ as $m,n\to\infty$, it holds that $p_{ji}\asymp\exp(\beta_{m,n})$. This paper studies in details how the divergence rate of $\beta_{m,n}$ affects the hypergraph structure as well as the corresponding estimation and inference procedures.
2.1 An initial study of $\beta_{m,n}$
In this section, we present a simple analysis of how the rate of $p_{ji}$ changes the hypergraph structure as an initial study of $\beta_{m,n}$. Let $\mathrm{Deg}(i):=\sum_{j\in[m]}\mathbf{1}_{\{i\in e_{j}\}}$ denote the number of hyperlinks that contain vertex $i$. We consider the following two concepts:
  • Null vertex: We call node $i$ a null vertex if $\mathrm{Deg}(i)=0$: if a vertex does not appear in any hyperlink in the observed hypergraph, we cannot learn anything about it.
  • Informative link: We call hyperlink $j$ an informative link if $|e_{j}|\neq 0$. When $|e_{j}|=0$, hyperlink $j$ does not provide any information for modeling the hypergraph nodes.
The following proposition shows that when $p_{ji}$’s are too small, an observed hypergraph generated from (1.1) contains null vertices and non-informative hyperlinks with high chance.
Proposition 2.1 (Necessary rate of $p_{ji}$).
In model (1.1), suppose $p_{ji}\leq\frac{a}{n}$ for all $j\in[m],i\in[n]$, and some constant $a>0$. Then $\lim_{m\to\infty}\lim_{n\to\infty}\mathbb{P}\big{(}\exists j\in[m],\text{ s.t.}|e_{j}|=0\mid\boldsymbol{\Xi}_{m,n}\big{)}=1$. By symmetry, suppose $p_{ji}\leq\frac{a}{m}$ for all $j\in[m],i\in[n]$, and some constant $a>0$ in (1.1). Then $\lim_{n\to\infty}\lim_{m\to\infty}\mathbb{P}\big{(}\exists i\in[n],\text{ s.t.}\mathrm{Deg}(i)=0\mid\boldsymbol{\Xi}_{m,n}\big{)}=1$.
The next proposition shows that the above rate of $p_{ji}$ is almost sufficient (up to a logarithmic factor) to ensure that there is no null vertex and non-informative hyperlink.
Proposition 2.2 (Sufficient rate of $p_{ji}$).
In model (1.1), suppose $p_{ji}\geq\frac{n^{\epsilon}}{n}$ for all $j\in[m],i\in[n]$ with $n^{\epsilon}=\Omega(\log m)$ for some $\epsilon\in(0,1)$ as $m,n\to\infty$. Then $\lim_{m,n\to\infty}\mathbb{P}\big{(}|e_{j}|\neq 0,~{}\forall j\in[m]\mid\boldsymbol{\Xi}_{m,n}\big{)}=1$. By symmetry, suppose $p_{ji}\geq\frac{m^{\epsilon}}{m}$ for all $j\in[m],i\in[n]$ with $m^{\epsilon}=\Omega(\log n)$ for some $\epsilon\in(0,1)$ as $m,n\to\infty$. Then $\lim_{m,n\to\infty}\mathbb{P}\big{(}\mathrm{Deg}(i)\neq 0,~{}\forall i\in[n]\mid\boldsymbol{\Xi}_{m,n}\big{)}=1$.
Recall that $p_{ji}\asymp\exp(\beta_{m,n})$ when $\beta_{m,n}\to-\infty$. Propositions 2.1 and 2.2 indicate the required rate of $\beta_{m,n}$ to ensure that a hypergraph avoids null vertices and non-informative links. The established rate is $\exp(\beta_{m,n})\asymp p_{ji}\gtrsim\epsilon_{m,n}/(m\wedge n)$ for some slowly diverging $\epsilon_{m,n}>0$ as $m,n\to\infty$. In Section 3, we explore the identification and estimation of parameters in $\boldsymbol{\Xi}_{m,n}$ and further investigate the rate of the order-adjusting parameter.
3 Identifiability, estimation and statistical theory
To streamline theoretical analysis and algorithmic implementation, we introduce a representation for hyperlinks using binary vectors. Specifically, for hyperlink $e_{j}$, we record $\mathbf{y}_{j}\in\{0,1\}^{n}$ where the $i$-th element $y_{ji}=\mathbf{1}_{\{i\in e_{j}\}}$. Denote $\boldsymbol{\alpha}=(\alpha_{1},\ldots,\alpha_{n})^{\top}\in\mathbb{R}^{n}$ and $\mathbf{Z}=(\mathbf{z}_{1},\ldots,\mathbf{z}_{n})^{\top}\in\mathbb{R}^{n\times K}$. Consider the logistic function applied component-wise to a vector $\mathbf{a}\in\mathbb{R}^{n}$ to yield $\sigma(\mathbf{a})=(\sigma(a_{i}),i\in[n])^{\top}\in[0,1]^{n}$. Then we define $\mathbf{p}_{j}=\sigma\big{(}\beta_{m,n}\mathbf{1}_{n}+\boldsymbol{\alpha}+\mathbf{Z}\mathbf{f}_{j}\big{)}$, and we have that $\{\mathbf{y}_{j}\}_{j\in[m]}$ are conditionally independent observations from multivariate Bernoulli distributions:
$\mathbf{y}_{j}\mid\boldsymbol{\Xi}_{m,n}\sim\mathrm{multiBernoulli}(\mathbf{p}_{j}),$
or equivalently, $y_{ji}$’s are conditionally independent Bernoulli random variables with means $p_{ji}=\mathbb{P}(y_{ji}=1\mid\boldsymbol{\Xi}_{m,n}).$ We use the superscript ``"``"``*"` ` ∗ " to denote the true model parameters: $\boldsymbol{\Xi}_{m,n}^{*}=\big{\{}\beta_{m,n}^{*},\{\alpha_{i}^{*}\}_{i=1}^{n},\{\mathbf{z}_{i}^{*}\}_{i=1}^{n},\{\mathbf{f}_{j}^{*}\}_{j=1}^{m}\big{\}}$. Denote $\boldsymbol{\alpha}^{*}=(\alpha^{*}_{1},\cdots,\alpha_{n}^{*})^{\top}\in\mathbb{R}^{n}$, $\mathbf{F}^{*}=(\mathbf{f}_{1}^{*},\cdots,\mathbf{f}_{m}^{*})^{\top}\in\mathbb{R}^{m\times K}$, $\mathbf{Z}^{*}=(\mathbf{z}_{1}^{*},\cdots,\mathbf{z}_{n}^{*})^{\top}\in\mathbb{R}^{n\times K}$, and $\boldsymbol{\Theta}^{*}=(\theta_{ji}^{*})_{m\times n}=\beta^{*}_{m,n}\mathbf{1}_{m}\mathbf{1}_{n}^{\top}+\mathbf{1}_{m}\boldsymbol{\alpha}^{*\top}+\mathbf{F}^{*}\mathbf{Z}^{*\top}\in\mathbb{R}^{m\times n}$. In the following, Section 3.1 considers the identifiability of $\beta^{*}_{m,n}$ and $\boldsymbol{\alpha}^{*}$. A constrained maximum likelihood estimator is studied and its $F$-consistency in the estimation for $\boldsymbol{\Theta}^{*}$, $\boldsymbol{\alpha}^{*}$ and $\beta^{*}_{m,n}$ is derived. Section 3.2 further considers the identification and estimation of all the individual components in $\boldsymbol{\Xi}_{m,n}^{*}$. Another constrained maximum likelihood estimator with additional constraints is studied. We show entry-wise consistency of the estimators and derive their asymptotic distributions. These results further enable consistent estimation and asymptotic inference for each $\theta^{*}_{ji}$ and $p_{ji}^{*}=\sigma(\theta^{*}_{ji})$.
3.1 $F$-consistent estimation for $\boldsymbol{\Theta}^{*}$, $\boldsymbol{\alpha}^{*}$ and $\beta_{m,n}^{*}$
Collect $\{\mathbf{y}_{j}\}_{j\in[m]}$ in a binary matrix such that ${\mathbf{Y}}=(\mathbf{y}_{1},\mathbf{y}_{2},\ldots,\mathbf{y}_{m})^{\top}\in\{0,1\}^{m\times n}$. The log-likelihood function is given by ${\cal L}(\beta,\boldsymbol{\alpha},\mathbf{F},\mathbf{Z}|{\mathbf{Y}}):=\sum_{j=1}^{m}\sum_{i=1}^{n}\big{\{}y_{ji}\theta_{ji}-\log\big{(}1+\exp(\theta_{ji})\big{)}\big{\}}$, with $\mathbf{F}=(\mathbf{f}_{1},\cdots,\mathbf{f}_{m})^{\top}$ and $\theta_{ji}=\beta+\alpha_{i}+\mathbf{f}_{j}^{\top}\mathbf{z}_{i}$. Consider the feasible region
[…]
$\max_{i,j}\{|\alpha_{i}|,\|\mathbf{z}_{i}\|_{2},\|\mathbf{f}_{j}\|_{2}\}\leq C_{1},-C_{\beta_{m,n}}\leq\beta\leq C_{2}\big{\}},$
where $C_{1},C_{2}>0$ are some constants, and $\mathbf{J}_{m}=\mathbf{I}_{m}-m^{-1}\mathbf{1}_{m}\mathbf{1}_{m}^{\top}$ is the centering operator. The boundedness constraint $\max_{i,j}\{|\alpha_{i}|,\|\mathbf{z}_{i}\|_{2},\|\mathbf{f}_{j}\|_{2}\}\leq C_{1}$ is similar to those employed in parameter estimation for fitting dyadic network latent space models (e.g., Ma et al., 2020). Note that we allow $\beta_{m,n}^{*}$ to diverge towards $\infty$- ∞ as $m,n\to\infty$, thus the parameter $C_{\beta_{m,n}}$ serves to control the rate of divergence of $\beta$ in the estimation, and the choice of the rate will be discussed in Remark 3.3 below.
We first examine the statistical property of the CMLE (constrained maximum likelihood estimator) defined as follows:
$(\hat{\beta},\hat{\boldsymbol{\alpha}},\hat{\mathbf{F}},\hat{\mathbf{Z}})=\max_{\beta,\boldsymbol{\alpha},\mathbf{F},\mathbf{Z}}$ ${\cal L}(\beta,\boldsymbol{\alpha},\mathbf{F},\mathbf{Z}|{\mathbf{Y}})~{}~{}\mathrm{s.t.}~{}~{}(\beta,\boldsymbol{\alpha},\mathbf{F},\mathbf{Z})\in{\cal F}_{1}(C_{\beta_{m,n}}).$ (3.2)
The core of (3.2) is to find a low-rank estimator for $\boldsymbol{\Theta}^{*}$ based on its logistic-transformed, noisy observations, and to recover $\beta_{m,n}^{*}$, $\boldsymbol{\alpha}^{*}$ from the estimate for $\boldsymbol{\Theta}^{*}$. For an estimator $\hat{\boldsymbol{\Theta}}\in\mathbb{R}^{m\times n}$ of $\boldsymbol{\Theta}^{*}$, we call $\hat{\boldsymbol{\Theta}}$ an $F$-consistent estimator if $(mn)^{-\frac{1}{2}}\|\hat{\boldsymbol{\Theta}}-\boldsymbol{\Theta}^{*}\|_{F}=o_{p}(1)$ as $m,n\to\infty$. Theorem 3.1 below shows the $F$-consistency of the CMLE (3.2).
Theorem 3.1 ($F$-consistent estimation of $\boldsymbol{\Theta}^{*}$).
Let $\hat{\boldsymbol{\Theta}}=\hat{\beta}\mathbf{1}_{m}\mathbf{1}_{n}^{\top}+\mathbf{1}_{m}\hat{\boldsymbol{\alpha}}^{\top}+\hat{\mathbf{F}}\hat{\mathbf{Z}}^{\top}$ where $(\hat{\beta},\hat{\boldsymbol{\alpha}},\hat{\mathbf{F}},\hat{\mathbf{Z}})$ is the solution to (3.2). For any $\epsilon>0$, as long as $\exp(\beta^{*}_{m,n})\gtrsim\log\big{(}(m\vee n)/\epsilon\big{)}/(m\wedge n)$ and $(\beta_{m,n}^{*},\boldsymbol{\alpha}^{*},\mathbf{F}^{*},\mathbf{Z}^{*})\in{\cal F}_{1}(C_{\beta_{m,n}})$, we have
$\|\hat{\boldsymbol{\Theta}}-\boldsymbol{\Theta}^{*}\|_{F}\lesssim\frac{(m\vee n)^{\frac{1}{2}}\exp(\beta^{*}_{m,n}/2)\big{\{}\log((m\vee n)/\epsilon\big{\}}^{\frac{1}{2}}}{\exp(-C_{\beta_{m,n}})},$ (3.3)
with probability at least $1-\epsilon$.
Remark 3.1.
The condition $\exp(\beta^{*}_{m,n})\gtrsim\log((m\vee n)/\epsilon)/(m\wedge n)$ in Theorem 3.1 suggests that as the number of vertices and the number of hyperlinks grow to infinity, the order of hyperlinks should be bounded below by a term proportional to $n\log((m\vee n)/\epsilon)/(m\wedge n)$. When $m\asymp n$, this order accords with the sufficient order in Proposition 2.2 and the necessary order in Proposition 2.1 up to a logarithmic factor.
Remark 3.2.
Theorem 3.1 also suggests that if $m\asymp n$, $\exp(\beta^{*}_{m,n})\asymp\exp(-C_{\beta_{m,n}})$, and $\exp(\beta^{*}_{m,n})\gg\log(m\vee n)/(m\wedge n)$, then the CMLE achieves $F$-consistent $\boldsymbol{\Theta}$ estimation, i.e., $(mn)^{-1/2}\|\hat{\boldsymbol{\Theta}}-\boldsymbol{\Theta}^{*}\|_{F}=o_{p}(1)$.
Remark 3.3 (A choice of $C_{\beta_{m,n}}$).
The choice of $C_{\beta_{m,n}}$ plays an important role in Theorem 3.1. As shown in (3.3), a larger $C_{\beta_{m,n}}$ leads to a worse estimation error bound. Meanwhile, $C_{\beta_{m,n}}$ needs to be large enough so that the condition $(\beta_{m,n}^{*},\alpha^{*},\mathbf{F}^{*},\mathbf{Z}^{*})\in{\cal F}_{1}(C_{\beta_{m,n}})$ in Theorem 3.1 could hold. Such a $C_{\beta_{m,n}}$ can be set in practice via $C_{\beta_{m,n}}=C^{\prime}\cdot\hat{C}_{m,n}$ where
$\hat{C}_{m,n}=-\log\Big{(}\frac{\sum_{j=1}^{m}\sum_{i=1}^{n}\mathbf{1}_{\{y_{ji}=1\}}}{mn}\Big{)}$
and $C^{\prime}\geq 1$ is a constant that does not depend on $m,n$. In this construction, $\hat{C}_{m,n}$ offers a reliable estimate of $\beta_{m,n}^{*}$ and the constant $C^{\prime}$ helps ensure that ${\cal F}_{1}(C_{\beta_{m,n}})$ contains the true parameters. The following proposition states the effectiveness of this construction.
Proposition 3.1.
Suppose that there exists $s_{m,n}\to\infty$ such that $\exp(\beta_{m,n}^{*})\gg s_{m,n}/(mn)$ as $m,n\to\infty$. Then the choice of $C_{\beta_{m,n}}$ in Remark 3.3 satisfies $C_{\beta_{m,n}}\asymp-\beta^{*}_{m,n}$ with probability approaching $1$ as $m,n\to\infty$. Consequently, there exist constants $C^{\prime}\geq 1$ in Remark 3.3 and $C_{1},C_{2}>0$ in (3.1) such that
$\mathbb{P}\big{\{}(\beta_{m,n}^{*},\boldsymbol{\alpha}^{*},\mathbf{F}^{*},\mathbf{Z}^{*})\in{\cal F}_{1}(C_{\beta_{m,n}})\big{\}}\to 1$
as $m,n\to\infty$.
We also note that the $F$-consistency result in Theorem 3.1 leads to the following corollary, which establishes $F$-consistency for the estimation of both ${\boldsymbol{\alpha}}^{*}$ and $\beta^{*}_{m,n}$.
Corollary 3.1 ($F$-consistent estimation of $\beta_{m,n}^{*}$ and $\boldsymbol{\alpha}^{*}$).
Let $(\hat{\beta},\hat{\boldsymbol{\alpha}},\hat{\mathbf{F}},\hat{\mathbf{Z}})$ be the solution to (3.2). For any $\epsilon>0$, as long as $\exp(\beta^{*}_{m,n})\gtrsim\log\big{(}(m\vee n)/\epsilon\big{)}/(m\wedge n)$ and $(\beta_{m,n}^{*},\boldsymbol{\alpha}^{*},\mathbf{F}^{*},\mathbf{Z}^{*})\in{\cal F}_{1}(C_{\beta_{m,n}})$, we have
$\|\hat{\boldsymbol{\alpha}}-\boldsymbol{\alpha}^{*}\|_{2}\lesssim\frac{(m\vee n)^{\frac{1}{2}}\exp(\beta^{*}_{m,n}/2)\big{\{}\log((m\vee n)/\epsilon\big{\}}^{\frac{1}{2}}}{m^{\frac{1}{2}}\exp(-C_{\beta_{m,n}})},$
$|\hat{\beta}-\beta_{m,n}^{*}|\lesssim\frac{(m\vee n)^{\frac{1}{2}}\exp(\beta^{*}_{m,n}/2)\big{\{}\log((m\vee n)/\epsilon\big{\}}^{\frac{1}{2}}}{m^{\frac{1}{2}}n^{\frac{1}{2}}\exp(-C_{\beta_{m,n}})},$
with probability at least $1-\epsilon$.
Similar to Remark 3.2, if $m\asymp n$, $\exp(\beta^{*}_{m,n})\asymp\exp(-C_{\beta_{m,n}})$, and $\exp(\beta^{*}_{m,n})\gg\log(m\vee n)/(m\wedge n)$, we obtain $F$-consistent estimation of both $\boldsymbol{\alpha}^{*}$ and $\beta^{*}_{m,n}$, i.e., $n^{-1/2}\|\hat{\boldsymbol{\alpha}}-\boldsymbol{\alpha}^{*}\|_{2}=o_{p}(1)$ and $|\hat{\beta}-\beta_{m,n}^{*}|=o_{p}(1)$.
3.2 Entry-wise consistent estimation and asymptotic inference for embedding vectors
In this section, we establish entry-wise consistent estimation and asymptotic inference for all the individual components in $\boldsymbol{\Xi}_{m,n}^{*}$. To make the embedding vectors identiafiable, further assumptions need to be made. The identification issue is due to the fact that for any full rank $\mathbf{A}\in\mathbb{R}^{K\times K}$, letting $\widetilde{\mathbf{f}}_{j}=\mathbf{A}^{\top}\mathbf{f}_{j}^{*}$ and $\widetilde{\mathbf{z}}_{i}=\mathbf{A}^{-1}\mathbf{z}_{i}^{*}$ leads to
$\widetilde{\mathbf{f}}_{j}^{\top}\widetilde{\mathbf{z}}_{i}=\mathbf{f}_{j}^{*\top}\mathbf{A}\mathbf{A}^{-1}\mathbf{z}^{*}_{i}={\mathbf{f}}_{j}^{*\top}\mathbf{z}_{i}^{*},$ (3.4)
which preserves the likelihood in (1.1), making individual embeddings not identifiable from observed data. To address the identifiability issue, we assume a set of embedding vector constraints commonly employed as follows:
[…]
The third constraint is from Section 2 for the identification of vertex degree parameters and the first two are for identifying the linear transformation $\mathbf{A}$. If the true embedding vectors do not satisfy the first two constraints, a transformation of the embeddings that satisfies the constraints can be performed as follows. Let ${\cal V}=\operatorname{diag}(\rho_{1}^{2},\cdots,\rho_{K}^{2})$ with $\rho_{1}^{2}>\rho_{2}^{2}>\cdots>\rho_{K}^{2}$ being the eigenvalues of $(mn)^{-1}(\mathbf{Z}^{*\top}\mathbf{Z}^{*})^{1/2}\mathbf{F}^{*\top}\mathbf{F}^{*}(\mathbf{Z}^{*\top}\mathbf{Z}^{*})^{1/2}$ and let $\boldsymbol{\Gamma}$ be a $K\times K$ matrix whose columns collect the corresponding eigenvectors. Let $\mathbf{G}=\big{(}\mathbf{Z}^{*\top}\mathbf{Z}^{*}/n\big{)}^{1/2}\boldsymbol{\Gamma}{\cal V}^{-\frac{1}{4}}$. The transformed $(\mathbf{Z}^{*},\mathbf{F}^{*})$ is $\big{(}\mathbf{Z}^{*}(\mathbf{G}^{\top})^{-1},\mathbf{F}^{*}\mathbf{G}\big{)}$. One interpretation of this transformation is that we fix $\mathbf{A}=\mathbf{G}$ in (3.4).
Another identifiability issue concerns with column sign flip. In particular, let $\mathbf{D}$ be a $K\times K$ diagonal matrix with $\{-1,1\}$ in its diagonal entries. Note that $\mathbf{D}$ is symmetric and $\mathbf{D}^{2}=\mathbf{I}$. Even when $\mathbf{F}^{*}$ and $\mathbf{Z}^{*}$ satisfy the above embedding vector constraints, if we let $\widetilde{\mathbf{F}}=\mathbf{F}^{*}\mathbf{D}$ and $\widetilde{\mathbf{Z}}=\mathbf{Z}^{*}\mathbf{D}$, $\widetilde{\mathbf{F}}$ and $\widetilde{\mathbf{Z}}$ would still satisfy the constraints and it holds that $\widetilde{\mathbf{F}}\widetilde{\mathbf{Z}}^{\top}=\mathbf{F}^{*}\mathbf{D}\mathbf{D}\mathbf{Z}^{*\top}=\mathbf{F}^{*}\mathbf{Z}^{*\top}$. Consequently, $\mathbf{F}^{*}$ and $\mathbf{Z}^{*}$ are identifiable up to column sign flips. We deal with this issue by fixing the sign of all coordinates in the first vertex embedding vector and adjust the estimators accordingly. In comparison to the identification of the linear transformation, the identification issue with column sign flipping is minor.
Under the above identifiability conditions, we establish entry-wise consistency and asymptotic normality in the estimation of all individual parameters in $\boldsymbol{\Xi}_{m,n}$. Note that due to the constraint $\sum_{i\in[n]}\alpha^{*}_{i}=0$, there are indeed $n$ rather than $n+1$ free parameters in $(\beta_{m,n}^{*},\boldsymbol{\alpha}^{*})$. For convenience in the theoretical analysis, we re-parameterize the model by integrating $\beta_{m,n}^{*}$ and $\boldsymbol{\alpha}^{*}$ into an $\boldsymbol{\alpha}^{\dagger}$ term, which results in un-centered vertex degree parameters. Specifically, we let $\alpha_{i}^{\dagger}=\beta_{m,n}^{*}+\alpha_{i}^{*}$ for $i\in[n]$. We now aim to estimate individual parameters in $(\boldsymbol{\alpha}^{\dagger},\mathbf{F}^{*},\mathbf{Z}^{*})$.
Let $\mathbf{f}=(\mathbf{f}_{1}^{\top},\ldots,\mathbf{f}_{m}^{\top})^{\top}$ and $\mathbf{z}=(\mathbf{z}_{1}^{\top},\ldots,\mathbf{z}_{n}^{\top})^{\top}$ denote vectorized versions of $\mathbf{F}$ and $\mathbf{Z}$, where $\mathbf{f}_{j}^{\top}\in\mathbb{R}^{K}$ is the $j$th row of $\mathbf{F}$ and $\mathbf{z}_{i}^{\top}\in\mathbb{R}^{K}$ is the $i$th row of $\mathbf{Z}$. Similarly, let $\mathbf{f}^{*}$ and $\mathbf{z}^{*}$ denote vectorized versions of $\mathbf{F}^{*}$ and $\mathbf{Z}^{*}$, respectively. Under the new parametrization, we rewrite the feasible region ${\cal F}_{1}(C_{\beta_{m,n}})$ in (3.1) in the following form:
[…]
$\|\boldsymbol{\alpha}-\sum_{i=1}^{n}\alpha_{i}/n\cdot\mathbf{1}_{n}\|_{\infty}\leq C_{4},~{}\max\{\max_{j\in[m]}\|\mathbf{f}_{j}\|_{2},\max_{i\in[n]}\|\mathbf{z}_{i}\|_{2}\}\leq C_{5},$
[…]
For the technical derivation of entry-wise consistent results and asymptotic distributions, we further set $C_{3}=-C_{3}^{\prime}\cdot C_{\beta_{m,n}}$ for some $C_{3}^{\prime}\in(0,1)$. As shown in the following proposition, this choice of $C_{3}$ retains the good properties of ${\cal F}_{1}(C_{\beta_{m,n}})$, where $C_{\beta_{m,n}}$ is taken as in Remark 3.3. The additional constraint for $C_{3}$ is mainly for the technical reason, and practically in algorithmic implementation, we find that the choice of $C_{3}$ can be flexible and does not significantly change the results.
Proposition 3.2.
Following the conditions in Proposition 3.1, there exist constants $C^{\prime}\geq 1$ in Remark 3.3 and $C_{3}^{\prime}\in(0,1),C_{4},C_{5}>0$ such that
$\mathbb{P}\big{\{}(\boldsymbol{\alpha}^{\dagger},\mathbf{f}^{*},\mathbf{z}^{*})\in{\cal F}_{2}(C_{\beta_{m,n}})\big{\}}\to 1$
as $m,n\to\infty$.
We also re-write the log-likelihood function as ${\cal L}({\mathbf{Y}}|\boldsymbol{\alpha},\mathbf{z},\mathbf{f})=\sum_{j=1}^{m}\sum_{i=1}^{n}\big{\{}y_{ji}(\alpha_{i}+\mathbf{f}_{j}^{\top}\mathbf{z}_{i})-\log\big{(}1+\exp(\alpha_{i}+\mathbf{f}^{\top}_{j}\mathbf{z}_{i})\big{)}\big{\}}$. We consider the following constrained maximum likelihood estimator
[…]
$\mathbf{J}_{m}\mathbf{F}=\mathbf{F},(\boldsymbol{\alpha},\mathbf{z},\mathbf{f})\in{\cal F}_{2}(C_{\beta_{m,n}}).$
Under the following conditions on the embedding vectors $\{\mathbf{f}_{j}^{*}\}_{j\in[m]}$ and $\{\mathbf{z}_{i}^{*}\}_{i\in[n]}$, the solution to (3.6) can achieve $F$-consistent estimation for $\{\boldsymbol{\alpha}^{\dagger},\mathbf{z}^{*},\mathbf{f}^{*}\}$. The $F$-consistency results will later be used for deriving entry-wise consistency results and asymptotic distributions.
Condition 3.1 (Eigen-structures and norm bounds of the embedding vectors).
We assume
$\rho_{\min}(m^{-1}\sum_{j=1}^{m}\mathbf{f}^{*}_{j}\mathbf{f}_{j}^{*\top})=\rho_{\min}(n^{-1}\sum_{i=1}^{n}\mathbf{z}^{*}_{i}\mathbf{z}_{i}^{*\top})\gtrsim 1,$ (3.7)
$\max_{j\in[m]}\|\mathbf{f}_{j}^{*}\|_{\infty}\lesssim 1$, and $\max_{i\in[n]}\|\mathbf{z}_{i}^{*}\|_{\infty}\lesssim 1$. Moreover, the eigenvalues of each of the $K\times K$ matrices $m^{-1}\sum_{j=1}^{m}\mathbf{f}^{*}_{j}\mathbf{f}^{*\top}_{j},n^{-1}\sum_{i=1}^{n}\mathbf{z}_{i}^{*}\mathbf{z}_{i}^{*\top}$ and $(m^{-1}\sum_{j=1}^{m}\mathbf{f}^{*}_{j}\mathbf{f}^{*\top}_{j})(n^{-1}\sum_{i=1}^{n}\mathbf{z}_{i}^{*}\mathbf{z}_{i}^{*\top})$ are distinct and their gaps are lower bounded by a constant.
Condition 3.1 ensures that the embedding vectors’ entries remain finite, same as the assumptions made in Section 2. Additionally, it requires non-degenerate second moment structures, thereby ensuring that each of the $K$ dimensions is non-trivial. The equal sign in (3.7) comes from the identifiability conditions. The distinctness assumption aims to ensure unique identification of the columns of $\mathbf{F}^{*}$ and $\mathbf{Z}^{*}$. Under these regular conditions, the $F$-consistency of the estimators in (3.6) are established as follows.
Proposition 3.3 ($F$-consistent estimation for $(\boldsymbol{\alpha}^{\dagger},\mathbf{z}^{*},\mathbf{f}^{*})$).
Let $(\hat{\boldsymbol{\alpha}},\hat{\mathbf{z}},\hat{\mathbf{f}})$ be the solution to (3.6). Under Condition 3.1, assume that $(\boldsymbol{\alpha}^{\dagger},\mathbf{z}^{*},\mathbf{f}^{*})\in{\cal F}_{2}(C_{\beta_{m,n}})$. Define
$\delta_{m,n}=\frac{(m\vee n)^{\frac{1}{2}}\exp(\beta_{m,n}^{*}/2)\big{\{}\log((m\vee n)\big{\}}^{\frac{1}{2}}}{m^{\frac{1}{2}}n^{\frac{1}{2}}\exp(-C_{\beta_{m,n}})}.$ (3.8)
Then as $(m,n)\to\infty$, we have
$n^{-\frac{1}{2}}\|\hat{\boldsymbol{\alpha}}-\boldsymbol{\alpha}^{\dagger}\|_{2}=O_{p}(\delta_{m,n}),~{}~{}n^{-\frac{1}{2}}\|\hat{\mathbf{z}}-\mathbf{z}^{*}\|_{2}=O_{p}(\delta_{m,n}),~{}~{}m^{-\frac{1}{2}}\|\hat{\mathbf{f}}-\mathbf{f}^{*}\|_{2}=O_{p}(\delta_{m,n}).$
Remark 3.4.
From Proposition 3.3, if $m\asymp n$, $\exp(\beta_{m,n}^{*})\asymp\exp(-C_{\beta_{m,n}})$, $\exp(\beta_{m,n}^{*})\gg\log(m\vee n)/(m\wedge n)$, we achieve the $F$-consistent estimation for $\boldsymbol{\alpha}^{\dagger}$, $\mathbf{z}^{*}$ and $\mathbf{f}^{*}$, i.e., $n^{-1/2}\|\hat{\boldsymbol{\alpha}}-\boldsymbol{\alpha}^{\dagger}\|_{2}=o_{p}(1)$, $(nK)^{-1/2}\|\hat{\mathbf{z}}-\mathbf{z}^{*}\|_{2}=o_{p}(1)$ and $(mK)^{-1/2}\|\hat{\mathbf{f}}-\mathbf{f}^{*}\|_{2}=o_{p}(1)$.
Next, we study entry-wise consistency properties and asymptotic normality of the estimators. This analysis presents several significant challenges. First, the minimum eigenvalue of the Hessian of the log-likelihood function is intricate to track due to identifiability issues and non-linearity of the problem. Moreover, the sparsity introduced by the diverging parameter $\beta_{m,n}^{*}$ as $m,n\to\infty$ further complicates the problem. To address these theoretical challenges, we introduce the following penalized and constrained maximum likelihood problem by using the method of Lagrange multipliers:
$\max_{\boldsymbol{\alpha},\mathbf{z},\mathbf{f}}{\cal Q}(\boldsymbol{\alpha},\mathbf{z},\mathbf{f}|\lambda_{m,n})$ $={\cal L}({\mathbf{Y}}|\boldsymbol{\alpha},\mathbf{z},\mathbf{f})-{\cal P}(\mathbf{z},\mathbf{f}|\lambda_{m,n}),\text{ s.t. }(\boldsymbol{\alpha},\mathbf{z},\mathbf{f})\in{\cal F}_{2}(C_{\beta_{m,n}}),$ (3.9)
with $\mbox{with }{\cal P}(\mathbf{z},\mathbf{f}|\lambda_{m,n})=$ $~{}\frac{\lambda_{m,n}mn}{8}\Big{\|}\text{diag}\Big{(}\frac{1}{n}\mathbf{Z}^{\top}\mathbf{Z}-\frac{1}{m}\mathbf{F}^{\top}\mathbf{F}\Big{)}\Big{\|}_{F}^{2}+\frac{\lambda_{m,n}mn}{2}\Big{\|}\frac{\mathbf{F}^{\top}\mathbf{1}_{m}}{m}\Big{\|}_{2}^{2}$
$~{}+\frac{\lambda_{m,n}mn}{2}\Big{\|}\text{ndiag}\Big{(}\frac{1}{m}\mathbf{F}^{\top}\mathbf{F}\Big{)}\Big{\|}_{F}^{2}+\frac{\lambda_{m,n}mn}{2}\Big{\|}\text{ndiag}\Big{(}\frac{1}{n}\mathbf{Z}^{\top}\mathbf{Z}\Big{)}\Big{\|}_{F}^{2},$
where $\lambda_{m,n}>0$ is the Lagrange multiplier parameter. For a squared matrix $\mathbf{A}$, $\text{diag}(\mathbf{A})$ stands for the diagonal matrix with diagonal entries equal to those of $\mathbf{A}$ and $\text{ndiag}(\mathbf{A})$ is the matrix containing upper triangular entries of $\mathbf{A}$, i.e., the $(i,j)$-th entry of $\text{ndiag}(\mathbf{A})$ is zero if $i\leq j$ and $A_{ij}$ if $i>j$. The penalty ${\cal P}(\mathbf{z},\mathbf{f}|\lambda_{m,n})$ in (3.9) is based on identifiability constraints in (3.6). The next proposition shows the equivalence between (3.9) and (3.6).
Proposition 3.4.
For any $\lambda_{m,n}>0$, the solutions to (3.6) and (3.9) are equivalent.
Since the solution remains unchanged for any $\lambda_{m,n}>0$, we can take advantage of the penalization term ${\cal P}(\mathbf{z},\mathbf{f}|\lambda_{m,n})$ in (3.9) when analyzing (3.6): in combination with the Hessian matrix of the log-likelihood function, the Hessian of the penalty function ${\cal P}$ expands the entire linear space $\mathbb{R}^{n(K+1)+mK}$ with high probability. Similar Lagrangian multiplier strategies have been used in the literature, such as Silvey (1959) and El-Helbawy & Hassan (1994) for Lagrangian multiplier test and Wang (2022) for generalized factor models with application to factor-augmented regressions. Due to space constraints, detailed theoretical analyses are in the supplementary material and we summarize the main results below. In Theorem 3.2, we establish entry-wise estimation error bounds for (3.6). We first introduce two conditions for achieving the entry-wise consistency results: one on the requirement for the order-adjusting parameter, and the other on the constraint parameters in (3.6).
Condition 3.2 (Hyperlink order).
There exists a small $\epsilon^{*}>0$ such that $\exp(\beta_{m,n}^{*})\gg(m\vee n)^{\epsilon^{*}}/(m\wedge n)$, as $m,n\to\infty$.
Condition 3.3 (Constraint parameter in (3.5)).
$C_{\beta_{m,n}}\asymp|\beta^{*}_{m,n}|$, as $m,n\to\infty$.
Condition 3.2 requires a minimum order for the hyperlinks which accords with the discussion in Section 2.1. Condition 3.3 indicates that the choice of $C_{\beta_{m,n}}$ should be reasonable based on the order-adjusting parameter.
Theorem 3.2 (Entry-wise consistent estimation of the embedding vectors and un-centered vertex degrees).
Let $(\hat{\boldsymbol{\alpha}},\hat{\mathbf{z}},\hat{\mathbf{f}})$ be the solution to (3.6). Under Conditions 3.1, 3.2 and 3.3, assume that $(\boldsymbol{\alpha}^{\dagger},\mathbf{z}^{*},\mathbf{f}^{*})\in{\cal F}_{2}(C_{\beta_{m,n}})$. Then we have
$\max\big{\{}\|\hat{\boldsymbol{\alpha}}-\boldsymbol{\alpha}^{\dagger}\|_{\infty},\|\hat{\mathbf{z}}-\mathbf{z}^{*}\|_{\infty},\|\hat{\mathbf{f}}-\mathbf{f}^{*}\|_{\infty}\big{\}}=O_{p}\Big{\{}\frac{(m+n)^{\epsilon}}{(m\wedge n)^{\frac{1}{2}}\exp(\beta_{m,n}^{*}/2)}\Big{\}}$
for any $\epsilon>0$, as $(m,n)\to\infty$.
Let $\mathbf{P}^{*}=(p^{*}_{ji})_{m\times n}=\sigma(\boldsymbol{\Theta}^{*})$. The next corollary demonstrates entry-wise consistent estimation of $\theta^{*}_{ji}$ and $p_{ji}^{*}$ based on (3.6).
Corollary 3.2 (Entry-wise consistent $\boldsymbol{\Theta}$ and $\mathbf{P}$ estimation).
Let $(\hat{\boldsymbol{\alpha}},\hat{\mathbf{z}},\hat{\mathbf{f}})$ be the solution to (3.6), $\hat{\boldsymbol{\Theta}}=(\hat{\theta}_{ji})_{m\times n}$ with $\hat{\theta}_{ji}=\hat{\alpha}_{i}+\hat{\mathbf{f}}_{j}^{\top}\hat{\mathbf{z}}_{i}$, and $\hat{\mathbf{P}}=\sigma(\hat{\boldsymbol{\Theta}})$. Then under the conditions in Theorem 3.2, for any $\epsilon>0$, we have
$\|\hat{\boldsymbol{\Theta}}-\boldsymbol{\Theta}^{*}\|_{\max}=O_{p}\Big{\{}\frac{(m+n)^{\epsilon}}{(m\wedge n)^{\frac{1}{2}}\exp(\beta_{m,n}^{*}/2)}\Big{\}}\text{ and }$
$\|\hat{\mathbf{P}}-\mathbf{P}^{*}\|_{\max}=O_{p}\Big{\{}\frac{(m+n)^{\epsilon}\exp(\beta_{m,n}^{*}/2)}{(m\wedge n)^{\frac{1}{2}}}\Big{\}}$
as $(m,n)\to\infty$.
In the following, we present asymptotic distributions for the estimators in (3.6). Define $\sigma^{\prime}(x)=\exp(x)/\{1+\exp(x)\}^{2}$ as the derivative of $\sigma(x)$. For ease of notation, we let $\mathbf{q}_{j}^{*}=(1,\mathbf{f}_{j}^{*\top})^{\top}$ and $\boldsymbol{\nu}_{i}=(\alpha_{i},\mathbf{z}_{i}^{\top})^{\top}$. Specify a pair $(j,i)\in[m]\times[n]$. We define
$\boldsymbol{\Sigma}_{(\boldsymbol{\nu}_{i},m,n)}=\bigg{(}\frac{1}{m}\sum_{j=1}^{m}\Big{[}\sigma^{\prime}(\theta^{*}_{ji})\exp(-\beta_{m,n}^{*})\mathbf{q}_{j}^{*}\mathbf{q}_{j}^{*\top}\Big{]}\bigg{)}^{-1}~{}\text{ and }$ (3.10)
$\boldsymbol{\Sigma}_{(\mathbf{f}_{j},m,n)}=\bigg{(}\frac{1}{n}\sum_{i=1}^{n}\Big{[}\sigma^{\prime}(\theta^{*}_{ji})\exp(-\beta_{m,n}^{*})\mathbf{z}_{i}^{*}\mathbf{z}_{i}^{*\top}\Big{]}\bigg{)}^{-1}.$
These definitions are strategically chosen to ensure that the magnitudes of $\boldsymbol{\Sigma}_{(\boldsymbol{\nu}_{i},m,n)}$ and $\boldsymbol{\Sigma}_{(\mathbf{f}_{j},m,n)}$ maintain a constant order as $m,n\to\infty$. This strategy is consistently applied in the remaining presentation of the asymptotic distribution results. The intention is to keep the “$\boldsymbol{\Sigma}$” terms and the ”$\mathrm{Var}$” terms that will be introduced later at a constant order, making it easier to see the rate of the asymptotic variance-covariance. Let $\boldsymbol{\nu}^{*}_{i}=(\alpha_{i}^{\dagger},\mathbf{z}_{i}^{*\top})^{\top}$ for $i\in[n]$. The following theorem presents asymptotic distributions for $(\hat{\alpha}_{i},\hat{\mathbf{z}}_{i}^{\top})^{\top}$ and $\hat{\mathbf{f}}_{j}$.
Theorem 3.3 (Individual asymptotic distributions).
Let $(\hat{\boldsymbol{\alpha}},\hat{\mathbf{z}},\hat{\mathbf{f}})$ be the solution to (3.6) and $\hat{\boldsymbol{\nu}}_{i}=(\hat{\alpha}_{i},\hat{\mathbf{z}}_{i}^{\top})^{\top}$. Under the conditions in Theorem 3.2, as $m,n\to\infty$, we have
[…]
Furthermore, let
${\boldsymbol{\Omega}}_{ji}^{(m,n)}=\begin{pmatrix}\frac{1}{m}\sum_{j=1}^{m}\Big{[}\sigma^{\prime}(\theta^{*}_{ji})\exp(-\beta_{m,n}^{*})\mathbf{q}_{j}^{*}\mathbf{q}_{j}^{*\top}\Big{]}&\frac{1}{m^{1/2}n^{1/2}}\sigma^{\prime}(\theta^{*}_{ji})\exp(-\beta_{m,n}^{*})\mathbf{q}_{j}^{*}\mathbf{z}_{i}^{*\top}\\ \frac{1}{m^{1/2}n^{1/2}}\sigma^{\prime}(\theta^{*}_{ji})\exp(-\beta_{m,n}^{*})\mathbf{z}_{i}^{*}\mathbf{q}_{j}^{*\top}&\frac{1}{n}\sum_{i=1}^{n}\Big{[}\sigma^{\prime}(\theta^{*}_{ji})\exp(-\beta_{m,n}^{*})\mathbf{z}_{i}^{*}\mathbf{z}_{i}^{*\top}\Big{]}\end{pmatrix},$
and define
$\boldsymbol{\Sigma}_{(\boldsymbol{\nu}_{i},\mathbf{f}_{j},m,n)}=\begin{bmatrix}\boldsymbol{\Sigma}_{(\boldsymbol{\nu}_{i},m,n)}&\mathbf{0}\\ \mathbf{0}&\boldsymbol{\Sigma}_{(\mathbf{f}_{j},m,n)}\end{bmatrix}{\boldsymbol{\Omega}}_{ji}^{(m,n)}\begin{bmatrix}\boldsymbol{\Sigma}_{(\boldsymbol{\nu}_{i},m,n)}&\mathbf{0}\\ \mathbf{0}&\boldsymbol{\Sigma}_{(\mathbf{f}_{j},m,n)}\end{bmatrix}.$
Then Theorem 3.4 obtains the joint asymptotic distribution for $(\hat{\alpha}_{i},\hat{\mathbf{z}}_{i},\hat{\mathbf{f}}_{j})$.
Theorem 3.4 (Joint asymptotic distributions).
Let $(\hat{\boldsymbol{\alpha}},\hat{\mathbf{z}},\hat{\mathbf{f}})$ be the solution to (3.6) and $\hat{\boldsymbol{\nu}}_{i}=(\hat{\alpha}_{i},\hat{\mathbf{z}}_{i}^{\top})^{\top}$. Under the conditions in Theorem 3.2, and further assume that $\exp(\beta_{m,n}^{*}/2)\gg(m^{\frac{1}{2}}+n^{\frac{1}{2}})(m+n)^{\epsilon^{*}}/(m\wedge n)$, then we have
[…]
as $m,n\to\infty$.
Note that as $m,n\to\infty$, $\boldsymbol{\Sigma}_{(\boldsymbol{\nu}_{i},\mathbf{f}_{j},m,n)}$ becomes block-wise diagonal, implying that $\hat{\boldsymbol{\nu}}_{i}$ and $\hat{\mathbf{f}_{j}}$ are asymptotically independent. Building upon the joint limit distributions in Theorem 3.4, we present limit distributions for estimates of $\theta_{ji}^{*}$ and $p_{ji}^{*}$ in Corollary 3.3.
Corollary 3.3 (Asymptotic distributions for $\theta_{ji}^{*}$ and $p_{ji}^{*}$ estimation).
Let
$\mathrm{Var}_{m,n}{(\theta_{ji})}=(m\wedge n)\begin{pmatrix}\mathbf{q}_{j}^{*}\\ \mathbf{z}_{i}^{*}\end{pmatrix}^{\top}\boldsymbol{\Sigma}_{(\boldsymbol{\nu}_{i},\mathbf{f}_{j},m,n)}\begin{pmatrix}\mathbf{q}_{j}^{*}\\ \mathbf{z}_{i}^{*}\end{pmatrix}~{}\text{ and}$
$\mathrm{Var}_{m,n}$ ${(p_{ji})}=(m\wedge n)(\sigma^{\prime\prime}(\theta_{ji}^{*})\exp(-\beta_{m,n}^{*}))^{2}\begin{pmatrix}\mathbf{q}_{j}^{*}\\ \mathbf{z}_{i}^{*}\end{pmatrix}^{\top}\boldsymbol{\Sigma}_{(\boldsymbol{\nu}_{i},\mathbf{f}_{j},m,n)}\begin{pmatrix}\mathbf{q}_{j}^{*}\\ \mathbf{z}_{i}^{*}\end{pmatrix}.$
Let $(\hat{\boldsymbol{\alpha}},\hat{\mathbf{z}},\hat{\mathbf{f}})$ be the solution to (3.6). Under the conditions in Theorem 3.4, let $\hat{\theta}_{ji}=\hat{\alpha}_{i}+\hat{\mathbf{f}}_{j}^{\top}\hat{\mathbf{z}}_{i}$ and $\hat{p}_{ji}=\sigma(\hat{\theta}_{ji})$, then we have
[…]
Theorem 3.3 and Corollary 3.3 provide the basis for constructing confidence intervals for $(\alpha_{i}^{\dagger},\mathbf{z}_{i}^{*},\mathbf{f}_{j}^{*})$, $\theta_{ji}^{*}$ and $p_{ji}^{*}$. To complete the inference task, we plug in the estimators $(\hat{\boldsymbol{\alpha}},\hat{\mathbf{z}},\hat{\mathbf{f}})$ into the covariance expressions. Note that $\beta_{m,n}^{*}$ does not need to be estimated and plugged in for inference implementation, as it cancels out in the covariance terms. We plug in $(\hat{\theta}_{ji},\hat{\mathbf{z}}_{i},\hat{\mathbf{f}}_{j})$ for $(\theta_{ji}^{*},\mathbf{z}_{i}^{*},\mathbf{f}_{j}^{*})$ in (3.10) and ${\boldsymbol{\Omega}}_{ji}^{(m,n)}$. Let $\widehat{\boldsymbol{\Sigma}}_{(\boldsymbol{\nu}_{i},m,n)}$ and $\widehat{\boldsymbol{\Sigma}}_{(\mathbf{f}_{j},m,n)}$ denote the plug-in estimators for ${\boldsymbol{\Sigma}}_{(\boldsymbol{\nu}_{i},m,n)}$ and ${\boldsymbol{\Sigma}}_{(\mathbf{f}_{j},m,n)}$, respectively. Let $\widehat{\mathrm{Var}}_{m,n}{(\theta_{ji})}$ and $\widehat{\mathrm{Var}}_{m,n}{(p_{ji})}$ be the plug-in estimators for ${\mathrm{Var}}_{m,n}{(\theta_{ji})}$ and ${\mathrm{Var}}_{m,n}{(p_{ji})}$, respectively. The following corollary establishes the consistency of the plug-in estimators.
Corollary 3.4.
Let $\delta^{\prime}_{m,n,\epsilon}={(m+n)^{\epsilon}}{(m\wedge n)^{-\frac{1}{2}}\exp(-\beta_{m,n}^{*}/2)}$. Under the conditions in Theorem 3.3, the plug-in estimators of the variance-covariance matrices satisfy that
  • $\|\widehat{\boldsymbol{\Sigma}}_{(\boldsymbol{\nu}_{j},m,n)}-\boldsymbol{\Sigma}_{(\boldsymbol{\nu}_{j},m,n)}\|_{F}=O_{p}(\delta^{\prime}_{m,n,\epsilon})$, $\|\widehat{\boldsymbol{\Sigma}}_{\mathbf{f}_{j},m,n}-\boldsymbol{\Sigma}_{\mathbf{f}_{j},m,n}\|_{F}=O_{p}(\delta^{\prime}_{m,n,\epsilon})$;
  • $|\widehat{\mathrm{Var}}_{m,n}{(\theta_{ji})}-{\mathrm{Var}}_{m,n}{(\theta_{ji})}|=O_{p}(\delta^{\prime}_{m,n,\epsilon})$, $|\widehat{\mathrm{Var}}_{m,n}{(p_{ji})}-{\mathrm{Var}}_{m,n}{(p_{ji})}|=O_{p}(\delta^{\prime}_{m,n,\epsilon})$,
for any $\epsilon>0$ as $m,n\to\infty$.
Combining Theorem 3.3, Corollaries 3.3 and 3.4 and applying Slutsky’s Theorem, we have that the asymptotic normality results hold for the plug-in variance-covariance.
4 Simulation studies
In this section, we conduct simulation studies to evaluate the performance of the proposed method and validate our theoretical findings. To compute the estimators, we employ a projected gradient descent algorithm (Ma et al., 2020) and use a universal singular value thresholding strategy (Chatterjee, 2015) for initialization. Details of the algorithms are in the supplementary materials. We consider various simulation setups where the embedding vectors are generated from a mixture of truncated multivariate Gaussian distributions. Specifically, when considering a $K$-dimensional latent space, we first randomly split the $n$ vertices into $K$ disjoint almost equal-sized groups, where “almost” means the sizes can differ by $1$ when $K$ does not divide $n$. Vertex embeddings are then independently generated as follows. For vertex $i$, if it is in the $k$th group, $\mathbf{z}^{*}_{i}$ would be generated from ${\cal N}_{[-1,1]}(\mathbf{e}_{k},\boldsymbol{\Sigma}_{\mathbf{z}})$, where $\mathbf{e}_{k}\in\{0,1\}^{K}$ is a column vector with the $k$th element being $1$ while other entries being 00, $\boldsymbol{\Sigma}_{\mathbf{z}}$ is the covariance matrix of the embedding vectors, and $[-1,1]$ indicates truncation around the mean by deviation 1 on each coordinate. The embedding vectors $\{\mathbf{f}^{*}_{j}\}_{j\in[m]}$ of the hyperlinks are centered i.i.d. realizations of ${\cal N}_{[-1,1]}(\mathbf{0},\boldsymbol{\Sigma}_{\mathbf{f}})$ so that $\sum_{j\in[m]}\mathbf{f}^{*}_{j}=\mathbf{0}$. For both $\boldsymbol{\Sigma}_{\mathbf{z}}$ and $\boldsymbol{\Sigma}_{\mathbf{f}}$, we adopt the autoregressive design where $\Sigma_{\mathbf{z},ij}=\Sigma_{\mathbf{f},ij}=0.2*\rho^{|i-j|}$ with $\rho\in\{0,0.5\}$. In particular, $\rho=0$ gives independent dimensions of the embedding vectors. The vertex degree heterogeneity parameters $\{\alpha_{i}^{*}\}\subset\mathbb{R}$ are i.i.d. realizations of $\mathrm{Uniform}[-1,1]$ and the order-adjusting parameter $\beta^{*}_{m,n}\in\mathbb{R}$ will be specified in each simulation setup. The hyperlinks $\mathbf{y}_{j}$’s are generated independently via $\mathbf{y}_{j}\sim\mathrm{multiBernoulli}\big{\{}\sigma(\boldsymbol{\theta}_{j}^{*})\big{\}}$, where $\boldsymbol{\theta}^{*}_{j}=\beta^{*}_{m,n}\mathbf{1}_{n}+\boldsymbol{\alpha}^{*}+\mathbf{Z}^{*}\mathbf{f}^{*}_{j}$. In Section 4.1, we investigate the effect of sample size, the order-adjusting parameter and the dimension of the latent space on the estimation of $\boldsymbol{\Theta}^{*}=(\boldsymbol{\theta}^{*}_{1},\ldots.\boldsymbol{\theta}^{*}_{m})^{\top}\in\mathbb{R}^{m\times n}$. In Section 4.2, we construct asymptotic confidence intervals for the individual parameters $(\boldsymbol{\alpha}^{\dagger},\mathbf{F}^{\mathbf{G}},\mathbf{Z}^{\mathbf{G}})$ and $(\boldsymbol{\Theta}^{*},\mathbf{P}^{*})$ based on the asymptotic normality results in Section 3.2, and evaluate their finite sample performance.
4.1 $\boldsymbol{\Theta}$ estimation
We first evaluate the effect of sample size and the dimension of the latent space together. We let $K\in\{2,3,4\}$, $\beta_{m,n}^{*}=-3$ and set $m=10n$ with $n\in\{100,200,\cdots,1000\}$. Figure 2 shows how the sample size changes the estimation error, with different setups of the latent space dimension. We can observe that as $(m,n)$ grows, the relative Frobenius error decreases proportionally to $(m\wedge n)^{-1/2}=n^{-1/2}$ in shape, which is consistent with the estimation error bound of Theorem 3.1. As the dimension of the latent embedding space grows, the estimation error grows correspondingly. Though not displayed in the error bound in Theorem 3.1 since we consider non-diverging latent space dimension in the theoretical analysis, an increase in $K$ would result in the increase of estimation error, due to the proof of Theorem 3.1 in the supplementary material. With higher correlations among dimensions of the embedding vectors, the estimation error remains almost the same and gets slightly higher as the dimension grows.
Refer to caption Refer to caption
(a) $\rho=0$ (b) $\rho=0.5$
Figure 2: Estimation error versus sample size based on $100$ Monte Carlo repetitions.
We further investigate the effect of the order-adjusting parameter under different latent space dimensions. We let $K\in\{2,3,4\}$, $m=10n=5,000$ and set $-\beta^{*}_{m,n}\in\{0.5,1,\cdots,3.5,4\}$. The results are displayed in Figure 3. We can see that as the value of $-\beta^{*}_{m,n}$ gets bigger, the estimation error grows proportional to $\exp(-\beta^{*}_{m,n})$. The effects of the dimension of the latent embedding space $K$ and the correlation structure among various dimensions are similar to those in Figure 2.
Refer to caption Refer to caption
(a) $\rho=0$ (b) $\rho=0.5$
Figure 3: Estimation error versus the order-adjusting parameter based on $100$ Monte Carlo repetitions.
Note that for conciseness, we focus on the presentation of the estimation error of $\boldsymbol{\Theta}^{*}$ in this subsection, while omitting the results for $\boldsymbol{\alpha}^{*}$ and $\beta_{m,n}^{*}$. The estimation errors for $\boldsymbol{\alpha}^{*}$ and $\beta_{m,n}^{*}$ are highly consistent with $\|\hat{\boldsymbol{\Theta}}-\boldsymbol{\Theta}^{*}\|_{F}$ by noticing that $\hat{\boldsymbol{\alpha}}-\boldsymbol{\alpha}^{*}=m^{-1}\mathbf{1}_{m}\mathbf{1}_{m}^{\top}(\hat{\boldsymbol{\Theta}}-\boldsymbol{\Theta}^{*})$ and $\hat{\beta}-\beta^{*}_{m,n}=n^{-1}\mathbf{1}_{n}^{\top}(\hat{\boldsymbol{\alpha}}-\boldsymbol{\alpha}^{*})$.
4.2 Confidence intervals
Refer to caption Refer to caption
Refer to caption Refer to caption
(a) $\beta^{*}_{m,n}=0$ (b) $\beta^{*}_{m,n}=-1$
Figure 4: Empirical coverage for $\{\alpha^{\dagger}_{i}\}_{i\in[n]}$, $\{Z_{ik}^{*}\}_{i\in[n],k\in[2]}$,$\{F_{jk}^{*}\}_{j\in[m],k\in[2]}$ and $\{\theta_{ii}^{*}\}_{i\in[m]}$, along with interval lengths. Each value comes from $1,000$ Monte Carlo repetitions.
Refer to caption Refer to caption
Figure 5: Empirical coverage and interval length for $\{p_{ii}^{*}\}_{i\in[m]}$. Each value comes from $1,000$ Monte Carlo repetitions.
In this section, we evaluate the finite sample performance of the constructed confidence intervals for our model parameters, including the embedding vectors, the un-centered degree heterogeneity parameters, $\theta_{ji}$’s, and $p_{ji}$’s. We set $K=2$ and vary $m=n\in\{400,800,1500,3000,6000\}$, $\beta^{*}_{m,n}\in\{0,-1\}$. After generating $(\beta^{*}_{m,n},\boldsymbol{\alpha}^{*},\mathbf{Z}^{*},\mathbf{F}^{*})$, $({\boldsymbol{\alpha}}^{\dagger},\mathbf{Z}^{*},\mathbf{F}^{*})$ are reconstructed accordingly. The $95\%$ confidence intervals are constructed following the methodologies outlined in Section 3.2. Figure 4 displays the empirical coverage values and confidence interval lengths for $(\boldsymbol{\alpha}^{\dagger},\mathbf{f}^{*},\mathbf{z}^{*},\boldsymbol{\Theta}^{*})$ and Figure 5 presents the results for $\mathbf{P}^{*}$, with both figures considering a range of $(m,n)$ values. They are presented in two separate plots due to the differing scales of $(\boldsymbol{\alpha}^{\dagger},\mathbf{f}^{*},\mathbf{z}^{*},\boldsymbol{\Theta}^{*})$ and $\mathbf{P}^{*}$, making it more informative to display their interval lengths separately. In Figure 4, the top box plots summarize empirical coverage values for each specified coverage target. For example, the “F” coverage target at $n=3000$ aggregates $6000$ empirical coverage values for $\{F_{jk}^{*}\}_{j\in[3000],k\in[2]}$. For the case where the coverage target is “Theta” or the probability term, the box plots only display empirical coverage values for $\theta_{ji}^{*}$ with $j=i$, to avoid overburdening the plot software with all $(j,i)$ pairs. The empirical coverage values are based on 1000 Monte Carlo simulations. The box plots of the confidence interval lengths are constructed similarly.
We can see that when $\beta^{*}_{m,n}=0$, the empirical coverage values for different coverage targets all approach the nominal $95\%$ level, as $m$ and $n$ increase. When $\beta^{*}_{m,n}=-1$, the empirical coverage values are not as favorable, highlighting the increased difficulty in making inference in the presence of sparsity. Furthermore, as $(m,n)$ increases, the lengths of the confidence intervals notably shrink. For fixed $\beta^{*}_{m,n}$, the length of the intervals appears proportional to $(m\wedge n)^{-1/2}$, which aligns well with our theory.
5 The co-citation hypergraph
In this section, we apply our latent embedding approach to a co-citation hypergraph. We use the Multi-Attribute Dataset on Statisticians (MADStat) collected and cleaned by Ji et al. (2022). MADStat consists of citation and authorship information of 83,331 published papers with 47,311 authors, spanning from 1975 to 2015 and covering 36 statistical journals. We focus on papers published in the 36 journals from 1991 to 2000 and consider the top $3,000$ authors by cited frequency within these years, which leaves us $14,984$ papers that cited these authors. In the co-citation hypergraph, each author is a node, and each paper forms a hyperlink: if an author is cited by a paper, they appear on the corresponding hyperlink. That is, for the $j$th hyperlink (paper) $\mathbf{y}_{j}\in\{0,1\}^{3000}$, $y_{ji}=1$ if the $i$th author is cited by the $j$th paper and $y_{ji}=0$ otherwise.
To obtain embedding vectors for the authors, we apply Algorithms 1 and 2 in the supplemental material to the $14,984$ hyperlinks of the $3,000$ authors, with the projection step based on Section 3.2. In Figure 6, we plot the embedding vectors along with $95\%$ confidence regions for a group of selected authors. The selected authors come from a joint set of authors with the highest estimated degree parameter values and a group of authors highlighted in the research map in Ji et al. (2022). The positioning of the embedding vectors shows relationship of recognized research interests, as evidenced by co-citation, among the authors between 1991 and 2000. For instance, the embedding vectors of Luke Tierney and Robert Kass are positioned almost identically, which indicates their possibly shared research interests from co-citation. Specifically, among the $14,984$ hyperlinks (papers), Tierney was cited in $349$ papers, while Kass received citations from $174$ papers, with $106$ of these papers citing both. Similarly, Larry Wasserman and Luis Pericchi have closely located embeddings with overlapping confidence regions. Wasserman was cited by 121 papers, Pericchi by 69, and they share 24 co-citations.
Figure 6: The author embedding vectors with $95\%$ confidence regions based on the co-citation hypergraph from 1991 to 2000.
Further, the sizes and orientations of the confidence regions could provide more insights on the hypergraph. From Theorem 3.3 and Corollary 3.4, the estimated covariance for author $i$’s embedding is $\big{\{}\big{(}\sum_{j=1}^{m}\sigma^{\prime}(\hat{\theta}_{ji})\hat{\mathbf{q}}_{j}\hat{\mathbf{q}}_{j}^{\top}\big{)}^{-1}\big{\}}_{2:3,2:3}$, where $\hat{\mathbf{q}}_{j}=(1,\hat{\mathbf{f}}_{j}^{\top})^{\top}$ for $j\in[m]$ and $\{\mathbf{A}\}_{2:3,2:3}$ denotes the sub-matrix of $\mathbf{A}$ with columns and rows indexed in $\{2,3\}$. First consider a simplified case with $\hat{\theta}_{ji}=\theta_{i}$ for all $j\in[m]$; then we have
$\big{(}\sum_{j=1}^{m}\sigma^{\prime}(\hat{\theta}_{ji})\hat{\mathbf{q}}_{j}\hat{\mathbf{q}}_{j}^{\top}\big{)}^{-1}=\big{\{}\sigma^{\prime}(\theta_{i})\big{\}}^{-1}\begin{bmatrix}m^{-1}&\mathbf{0}_{2}^{\top}\\ \mathbf{0}_{2}&\big{(}\sum_{j=1}^{m}\hat{\mathbf{f}}_{j}\hat{\mathbf{f}}_{j}^{\top}\big{)}^{-1}\end{bmatrix}$
due to the fact that $\sum_{j=1}^{m}\hat{\mathbf{f}}_{j}=\mathbf{0}_{2}$. Consequently, the estimated covariance for author $i$’s embedding would be $\boldsymbol{\Sigma}_{i}:=\big{\{}\sigma^{\prime}(\theta_{i})\big{\}}^{-1}\big{(}\sum_{j=1}^{m}\hat{\mathbf{f}}_{j}\hat{\mathbf{f}}_{j}^{\top}\big{)}^{-1}$. The magnitude of $\boldsymbol{\Sigma}_{i}$ is determined by $m^{-1}$ times the inverse of the magnitude of the “sample covariance” quantity $m^{-1}\sum_{j=1}^{m}\hat{\mathbf{f}}_{j}\hat{\mathbf{f}}_{j}^{\top}$ and the magnitude of $\{\sigma^{\prime}(\theta_{i})\}^{-1}$. The orientation of $\boldsymbol{\Sigma}_{i}$ is the orientation of the ellipse of the “sample covariance” $m^{-1}\sum_{j=1}^{m}\hat{\mathbf{f}}_{j}\hat{\mathbf{f}}_{j}^{\top}$, with major/minor axes being flipped. This is due to the fact that $m^{-1}\sum_{j=1}^{m}\hat{\mathbf{f}}_{j}\hat{\mathbf{f}}_{j}^{\top}$ shares the same eigen-space as $(m^{-1}\sum_{j=1}^{m}\hat{\mathbf{f}}_{j}\hat{\mathbf{f}}_{j}^{\top})^{-1}$ while the eigen-values are inversed.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 7: Zoom-in versions of the embeddings in Figure 6 for the four authors discussed and corresponding paper embeddings.
In practical scenarios where $\hat{\theta}_{ji}$ vary across different $j$, the estimated covariance matrices are the inverses of weighted second moments of $\hat{\mathbf{q}}_{j}$’s, the analytic forms of which are rather challenging to evaluate directly. We use Figure 7 to explore the relationship between the shape of the confidence regions and the paper embeddings. At the top of Figure 7, we zoom in two parts in Figure 6. Each of these two parts highlights two author embeddings along with their confidence regions in the plots while omitting other authors. Beneath each zoomed-in part, we display the paper embeddings corresponding to the two authors. We then compare the sizes and orientations of the confidence regions for two pairs of authors: Tierney and Kass, and Pericchi and Wasserman, respectively. Despite the similarity in the magnitude and distribution of paper embeddings for Tierney and Kass, there is a notable difference in the size of their confidence regions. This is due to their different scales of $\sigma^{\prime}(\hat{\theta}_{ji})$’s, which accord with different numbers of citations they received as discussed before. The median value of $\sigma^{\prime}(\hat{\theta}_{ji})$ for Tierney is $1.57\times 10^{-3}$, while the median value for Kass is $6.90\times 10^{-4}$. Such a difference in the scale of $\sigma^{\prime}(\hat{\theta}_{ji})$’s leads to a larger confidence region for Kass’s embedding. Wasserman and Pericchi have confidence regions of similar sizes, while their orientations are different. This can be explained by their corresponding paper embeddings. The majority of their paper embeddings are closely located and similarly distributed. However, a few paper embeddings at the bottom of the plot (indicated by red rectangulars) drift the general distribution of the embeddings of papers citing Wasserman counterclockwise, resulting in a slight rotation in his author embedding’s confidence region. These observations shed some light on more understanding of the interactions among the vertices (authors) and their connections through individual hyperlinks.
In addition, another interesting observation is that the embeddings of Trevor Hastie and Robert Tibshirani in Figure 6, despite their frequent collaborations, do not locate as close as one might expect. We zoom-in their author embeddings along with corresponding confidence regions and plot the embeddings of papers citing their work between 1991 and 2000 in Figure 8. Notably, a subset of these papers, particularly those positioned in the upper right part of the plot (marked via red rectangles), only cite Tibshirani but not Hastie. These papers all reference “Efron, B. and Tibshirani, R. J. (1986) Bootstrap methods for standard errors, confidence intervals and other measures of statistical accuracy (with discussion). Statistical Science $\mathbf{1}$(1):54-75” (Efron & Tibshirani, 1986), which is extensively cited in the data set. Interestingly, in the data set, none of the papers co-authored by Hastie have “bootstrap” in their titles, whereas this term appears multiple times in Tibshirani’s co-authored papers. These observations unveil a difference in terms of research interests between Hastie and Tibshirani according to their co-citation information between 1991-2000, helping explain the different positioning of Hastie’s and Tibshirani’s embeddings. Also notably, Tibshirani’s embedding has a smaller confidence region compared to Hastie’s. This is again due to different scales of their $\sigma^{\prime}(\hat{\theta}_{ji})$’s. Specifically, the median value of $\sigma^{\prime}(\hat{\theta}_{ji})$ for Hastie is $2.26\times 10^{-3}$, in contrast to Tibshirani’s which is $1.09\times 10^{-2}$. The different scales of ${\sigma}^{\prime}(\hat{\theta}_{ji})$’s accord with their different frequencies of getting cited in the data set: Hastie by $283$ papers and Tibshirani by $347$.
Refer to caption Refer to caption
Figure 8: Author embeddings of Hastie T. and Thishirabi, R. and embeddings of papers citing them.
6 Discussion
This paper introduces a novel latent embedding approach for modeling high-dimensional general hypergraphs. The core objective of this methodology is to address pervasive and intricate challenges in hypergraph modeling: non-uniformity and multiplicity of hyperlinks. While we consider embedding the network vertices and the hyperlinks in the Euclidean space in this paper, the latent embedding approach can serve as a general framework for modeling hypergraphs, with various choices of latent spaces, similarity measures, and activation functions. Under the Euclidean inner product considered in this paper, the embedding vectors from our methodology provide meaningful interpretations. Specifically, the likelihood of any particular group of vertices appearing on a certain hyperlink is determined by the similarity between the hyperlink’s embedding vector and the centroid of the vertices’ embedding vectors. To handle the complex nature of real-world hypergraphs, we further integrate degree heterogeneity parameters, which serve to adjust the vertices’ different popularity, and include an order-adjusting parameter to model the overall order of the hyperlinks. The interpretability of our approach builds upon the integration of these parameters. Identifiability, estimation, and asymptotic behaviors are then studied within this framework.
In our theoretical analysis, we utilize the constraint sets in the estimation procedure and study a corresponding penalized maximum likelihood estimator (PMLE). Demonstrating the equivalence between the constrained maximum likelihood estimator (CMLE) and the PMLE, the construction of the PMLE facilitates a more accurate characterization of the CMLE’s properties, which yields the results in entry-wise consistency and asymptotic normality in this paper. In general, this Lagrangian multiplier type strategy has the potential for broad applications in a variety of constrained statistical estimation and inference problems. Moreover, our theoretical analysis studies the impact of the order-adjusting parameter on both estimation and inference procedures. We show that the exponential of the order-adjusting parameter is proportional to the expected order of the hyperlinks divided by the number of vertices. We require that the expected order of hyperlinks should be at least proportional to $n\cdot\log(m\vee n)/(m\wedge n)$ for $F$-consistent estimation and to $n(m\vee n)^{\epsilon^{*}}/(m\wedge n)$ for some $\epsilon^{*}>0$ for entry-wise consistent estimation. For valid inference, we further need the expected order to be at least proportional to $n(m+n)^{1+\epsilon^{*}}/(m\wedge n)^{2}$. These technical results are not only of theoretical interest, but also useful to guide practitioners on how and when to apply the latent embedding approach. By understanding the conditions necessary for the effective application of the latent embedding approach, users are better equipped to leverage this methodology in appropriate contexts.
Building on the inference results of this work, various downstream tasks in hypergraph modeling, such as hyperlink prediction and community detection, can be handled. We leave detailed explorations on these downstream tasks for future research. Another important future work concerns variations of the latent embedding approach beyond the linear inner product framework. For instance, Ma et al. (2020) studied a general model class adhering to the Schoenberg condition (Schoenberg, 1937, 1938), which goes beyond the linear inner product-based latent space model for dyadic networks. While the utilization of linear inner product in (2.1) provides a clear and intuitive interpretation, there is significant interest in developing and assessing embedding methods using alternative inner product operations. Comparison between these operations with the linear inner product could bring new insights. Further appealing research avenues include investigating the choice of the link (activation) function and determining the optimal latent space dimension $K$. The selection of appropriate modeling approaches is also of great interest, for instance, adopting the study in Li et al. (2020).